OpenStudy (anonymous):

please help me!!! solve each system of equations using elimination and/or substitution. express your answer as an ordered pair. 5x-2y=3 2x+7y=9

OpenStudy (anonymous):

Ok first you have to pick an equation and start isolating either the "x" or "y" variables.

OpenStudy (anonymous):

okay, so like... 2x=9-7y?

OpenStudy (anonymous):

Unless you prefer to do the elimination method

OpenStudy (anonymous):

i think i'll stick to substitution. don't wanna confuse the situation more!

OpenStudy (anonymous):

ok so you decide to isolate the "x" so now you would divde everything by 2 on both sides

OpenStudy (anonymous):

Since you would just get a big old decimal just leave it as a fraction

OpenStudy (anonymous):

x=9/2 - 7/2y

OpenStudy (anonymous):

okay......... x= 4.5-3.5

OpenStudy (anonymous):

ohhhhhhhkkkkkkaaaaaayyyyy... wow that is a lot easier the way you put it:)

OpenStudy (anonymous):

Actually if it end there you could use the decimals. I thought it would go on forever

OpenStudy (anonymous):

yeh, division sucks like that:)

OpenStudy (anonymous):

You can use either way

OpenStudy (anonymous):

anyway now "technically" we know that x = 4.5-3.5

OpenStudy (anonymous):

so now we would substitute all the "x" values for that small equation

OpenStudy (anonymous):

alright so like do 5(4.5)? hold up before i continue am i doing this right???

OpenStudy (anonymous):

yeah k think i got it... 5(4.5) - 2(3.5)=3?

OpenStudy (anonymous):

I didn't see your post correct like that, but you have to include the entire equation cause "x" doesn't just equal that

OpenStudy (anonymous):

"x" equals 4.5 - 3.5

OpenStudy (anonymous):

and don't forget the "y" at the end

OpenStudy (anonymous):

so x actually equals 4.5 - 3.5y?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

so rewrite the equation

OpenStudy (anonymous):

okay- 5(4.5-3.5y) wait what about the y do i just put 3.5?

OpenStudy (anonymous):

at the end?

OpenStudy (anonymous):

no add the "y" that's part of x's value

OpenStudy (anonymous):

so it would be: 5(4.5-3.5y) -2y = 3

OpenStudy (anonymous):

k- so 5(4.5 - 3.5y) -2y = 3 22.5-17.5) - 2y=3?

OpenStudy (anonymous):

Almost the "y" doesn't dissapear

OpenStudy (anonymous):

k so 22.5- 17.5Y-2y=3?

OpenStudy (anonymous):

Now we would add like terms so that we only have one y

OpenStudy (anonymous):

so -17.5 -2y

OpenStudy (anonymous):

-17.5y*

OpenStudy (anonymous):

forgot the "y"

OpenStudy (anonymous):

22.5-15.5y=3?

OpenStudy (anonymous):

No! This is a really bad mistake in math remember that a negative plus a negative is just like adding so it would be: -19.5y

OpenStudy (anonymous):

so.... 22.5-19.5y=3?

OpenStudy (anonymous):

yes and now just treat it like a normal equation

OpenStudy (anonymous):

wait but doesn't 22.5-19.5=3?

OpenStudy (anonymous):

i mean as you can see there is definately a reason i am here so maybe i'm just dumb

OpenStudy (anonymous):

so? our goal is to find out what the heck "y" is so we would use the opposite of the order of operations

OpenStudy (anonymous):

so 3 - 22.5 = ?

OpenStudy (anonymous):

-19.5y?

OpenStudy (anonymous):

no neither of them have "y" so you can't just make another "y" magicly appear

OpenStudy (anonymous):

well that just means it equals 19.5 correct?

OpenStudy (anonymous):

-19.5 so now we have: -19.5y = -19.5 any ideas what to do?

OpenStudy (anonymous):

it equals 1 cuz they're the same or something right?

OpenStudy (anonymous):

YES a negative divide by a negative is a positive

OpenStudy (anonymous):

so now we know that y = 1

OpenStudy (anonymous):

But what da heck does "x" equal?

OpenStudy (anonymous):

k so like i said 1 in the end the entire answer would be (1,1)

OpenStudy (anonymous):

Why would you say that? have you found "x" already?

OpenStudy (anonymous):

it equals 1 cuz they're the same or something right?

OpenStudy (anonymous):

well that just means it equals 19.5 correct?

OpenStudy (anonymous):

No, now what were gonna do is find "x" the way we did in the beggining

OpenStudy (anonymous):

So pick your favorite equation

OpenStudy (anonymous):

summer assignment... answer sheet... need to know how to show my work... all that good jazz.... so i came on here so someone could help me better understand how to do it and i completely understand now.. thank you... you are my math hero, etc.... :)

OpenStudy (anonymous):

But you know how to do the last step right?

OpenStudy (anonymous):

yes, don't worry thank you so so so so much;)

OpenStudy (anonymous):

:) Just in case what you do is grab one of the equations and substitute the "y" for 1 and just solve so yes it would be (1, 1)

OpenStudy (anonymous):

thank you very much:)